# Computational Hypergraph Discovery, a Gaussian Process framework for connecting the dots

### v1

@@ Abstract.
Most scientific challenges can be framed into one of the following three levels of complexity of function approximation. Type 1: Approximate an unknown function given input/output data. Type 2: Consider a collection of variables and functions, some of which are unknown, indexed by the nodes and hyperedges of a hypergraph (a generalized graph where edges can connect more than two vertices). Given partial observations of the variables of the hypergraph (satisfying the functional dependencies imposed by its structure), approximate all the unobserved variables and unknown functions. Type 3: Expanding on Type 2, if the hypergraph structure itself is unknown, use partial observations of the variables of the hypergraph to discover its structure and approximate its unknown functions. While most Computational Science and Engineering and Scientific Machine Learning challenges can be framed as Type 1 and Type 2 problems, many scientific problems can only be categorized as Type 3. Despite their prevalence, these Type 3 challenges have been largely overlooked due to their inherent complexity. Although Gaussian Process (GP) methods are sometimes perceived as well-founded but old technology limited to Type 1 curve fitting, their scope has recently been expanded to Type 2 problems. In this paper, we introduce an interpretable GP framework for Type 3 problems, targeting the data-driven discovery and completion of computational hypergraphs. Our approach is based on a kernel generalization of (1) Row Echelon Form reduction from linear systems to nonlinear ones and (2) variance-based analysis. Here, variables are linked via GPs, and those contributing to the highest data variance unveil the hypergraph’s structure. We illustrate the scope and efficiency of the proposed approach with applications to (algebraic) equation discovery, network discovery (gene pathways, chemical, and mechanical), and raw data analysis.
1. Introduction.
1.1. The three levels of complexity of function approximation.
Alfred North Whitehead stated in 1911: "Civilization advances by extending the number of important operations we can perform without thinking about them." The automation of arithmetic and calculus through the introduction of calculators and computers serves as a testament to such transformative shifts. In line with this perspective, the resolution of most scientific challenges could be facilitated by automating the resolution of the following three levels of increasing complexity of function approximation (see Fig. 1).
  • Type 1 (Regression): Approximate an unknown function given (possibly noisy) input/output data.
  • Type 2 (Hypergraph Completion): Consider a collection of variables and functions, some of which are unknown, indexed by the nodes and hyperedges of a hypergraph (a generalized graph where edges can connect more than two vertices). Given partial observations of the variables of the hypergraph (satisfying the functional dependencies imposed by its structure), approximate all the unobserved variables and unknown functions.
  • Type 3 (Hypergraph Discovery): Expanding on Type 2, if the hypergraph structure itself is unknown, use partial observations (of subsets of the variables of the hypergraph) to discover its structure (the hyperedges and possibly the missing vertices) and then approximate its unknown functions and unobserved variables.
@@
The scope of these problems is immense. Numerical approximation, Supervised Learning, and Operator Learning can all be formulated as type 1 problems (possibly with functional inputs/outputs spaces). Type 2 problems include (see Fig. 4) solving and learning (possibly stochastic) ordinary or partial differential equations, Deep Learning, dimension reduction, reduced-ordered modeling, system identification, closure modeling, etc. The scope of Type 3 problems extends well beyond Type 2 problems and includes applications involving model/equation/network discovery and reasoning with raw data. While most problems in Computational Sciences and Engineering (CSE) and Scientific Machine Learning (SciML) can be framed as Type 1 and Type 2 challenges, many problems in science can only be categorized as Type 3 problems. Despite their prevalence, these Type 3 challenges have been largely overlooked due to their inherent complexity.
Causal inference methods and probabilistic graphs, as discussed in Section 3.3, and sparse regression methods [11, 4], offer potential avenues for addressing Type 3 problems. However, it is important to note that their application to these problems necessitates additional assumptions. Causal inference models, for instance, typically assume randomized data and some level of access to the data generation process or its underlying distributions. We emphasize that our proposed approach does not seek to replace causal inference methods but rather seeks to encode a different type of information into the structure of the graph, i.e., the functional dependencies between variables (see Sec. 3.3 for a more detailed discussion). Sparse regression methods, on the other hand, rely on the assumption that functional dependencies have a sparse representation within a known basis. In this paper, we do not impose these assumptions, and thus, these particular techniques may not be applicable.
Figure 2. Formal description of Type 2 problems.
1.2. Type 2 problems: Formal description and GP-based Computational Graph Completion
1.2.1. Formal description of Type 2 problems
1.2.2. Reminder on Computational Graph Completion for Type 2 problems
Although Gaussian Process (GP) methods are sometimes perceived as a well-founded but old technology limited to curve fitting (Type 1 problems), they have recently been generalized, beyond Type 1 problems [26, 27, 37, 35, 16, 17, 36, 3], to an interpretable framework (Computational Graph Completion or CGC [28]) for solving Type 2 problems [29, 5, 2, 6, 8, 18], all while maintaining the simple and transparent theoretical and computational guarantees of kernel/optimal recovery methods [22, 26].
Within the context of Sec. 1.2.1, the proposed GP solution to Type 2 problems is to simply replace unknown functions by GPs and compute their Maximum A Posteriori (MAP)/Maximum Likelihood Estimation (MLE) estimators given available data and constraints imposed by the structure of the graph. Taking into account the example depicted in Fig. 2, and substituting $f_{5,1}$, $f_{1,2}$, and $f_{3,6}$ with independent GPs, each with kernels $K,G$, and $\Gamma$ respectively, the objective of this MAP solution becomes minimizing $\|f_{5,1}\|_{K}^{2}+\|f_{1,2}\|_{G}^{2}+\|f_{3,6}\|_{\Gamma}^{2}$ (writing $\|f\|_{K}$ for the RKHS norm of $f$ induced by the kernel $K$) subject to the constraints imposed by the data and the functional dependencies encoded into the structure of the graph.
1.2.3. A system identification example.
In order to exemplify Computational Graphical Completion (CGC), consider the system identification problem depicted in Fig.3, sourced from [28]. Our objective is to identify a nonlinear electric circuit, as illustrated in Fig.3.(a), from scarce measurement data. The nonlinearity of the circuit emanates from the resistance, capacitance, and inductances, which are nonlinear functions of currents and voltages, as shown in Fig. 3.(b). Assuming these functions to be unknown, along with all currents and voltages as unknown time-dependent functions, we operate the circuit between times 00 and $10$. Measurements of a subset of variables, representing the system’s state, are taken at times $t_{s}=s/10$ for $s\in{0,\ldots,99}$. Given these measurements, the challenge arises in approximating all unknown functions that define currents and voltages as time functions, capacitance as a voltage function, and inductances and resistance as current functions. Fig. 3.(c) displays the available measurements, which are notably sparse, preventing us from reconstructing the underlying unknown functions independently. Thus, their interdependencies must be utilized for approximation. It is crucial to note that the system’s state variables are interconnected through functional relations, as per Kirchhoff’s laws for this nonlinear electric circuit, illustrated in Fig.3.(d). These functional dependencies can be conceptualized as a computational graph, depicted in Fig.3.(e), where nodes represent variables and directed edges represent functions. Known functions are colored in black, unknown functions in red, and round nodes aggregate variables, meaning edges map groups of variables, forming a hypergraph. The CGC solution involves substituting the graph’s unknown functions with Gaussian Processes (GPs), which may be independent or correlated, and then approximating the unknown functions with their Maximum A Posteriori (MAP) estimators, given the available data and the functional dependencies embedded in the graph’s structure. Fig. 3.(f) showcases the true and recovered functions, demonstrating a notably accurate approximation despite the data’s scarcity.
This simple example generalizes to an abstract framework detailed in [28]. This framework has a wide range of applications because most problems in CSE can also be formulated as completing computational graphs representing dependencies between functions and variables, and they can be solved in a similar manner by replacing unknown functions with GPs and by computing their MAP/EB estimator given the data. These problems include (see Fig. 4) solving and learning PDEs [5, 2, 6], learning SDEs [8], learning dynamical systems from partial observations [18], Deep Learning [29], etc. Although ANNs are oftentimes used to solve Type 1 problems, their training (the identification of unknown functions encoded into their layers) is a Type 2 problem as illustrated in Fig. 4 for a ResNet (see [29] for details on the proposed GP solution/regularization).
Figure 4. Examples of problems that can be solved as Type 2 problems.
1.3. Examples of Computational Hypergraph Discovery (Type 3) problems
While most problems in Computer Science and Engineering can be framed and resolved as computational graph or hypergraph completion tasks, some scientific challenges necessitate first uncovering the hypergraph’s structure before completing it. To illustrate the need for our proposed Computational Hypergraph Discovery framework, we will describe examples of these challenges, some of which are addressed in this paper.
1.3.1. Recovery of chemical reaction networks.
One such problem is the recovery of chemical reaction networks from concentration snapshots. As an example consider the hydrogenation of ethylene $(\text{C}_{2}\text{H}_{4})$ into ethane $(\text{C}_{2}\text{H}_{6})$ (described in [19]). A proposed mechanism for this chemical reaction, represented by
[…]
C2H4 + H2 -> C2H6
is
[…]
[math-space=0pt]H2 <=>[ $k_{1}$ ][ $k_{-1}$ ] 2 H
C2H4 + H ->[ $k_{2}$ ] C2H5
C2H5 + H ->[ $k_{3}$ ] C2H6
The problem is that of recovering the underlying chemical reaction network from snapshots (illustrated in Fig. 5) of concentrations $[H_{2}]$, $[H]$, $[C_{2}H_{4}]$ and $[C_{2}H_{5}]$ and their time derivatives. $\frac{d[H_{2}]}{dt}$, $\frac{d[H]}{dt}$, $\frac{d[C_{2}H_{4}]}{dt}$ and $\frac{d[C_{2}H_{5}]}{dt}$.
Figure 5. Pairwise scatter plot of the derivative and non-derivative variables of the chemistry problem
1.3.2. Data analysis.
Another problem is data analysis. As an example, consider the analysis of COVID-19 data from Google [^1] The dataset can be accessed here. Focus on a single country, France, to ensure consistency in the data and avoid considering cross-border variations that are not directly reflected in the data. Select 31 variables that describe the state of the country during the pandemic, spanning over 500 data points, with each data point corresponding to a single day. These variables are categorized as the following datasets:
  • Epidemiology dataset: Includes quantities such as new infections, cumulative deaths, etc.
  • Hospital dataset: Provides information on the number of admitted patients, patients in intensive care, etc.
  • Vaccine dataset: Indicates the number of vaccinated individuals, etc.
  • Policy dataset: Consists of indicators related to government responses, such as school closures or lockdown measures, etc.
Some of these variables are illustrated in Fig. 6. The problem is then to analyze this data and identify possible hidden functional relations between these variables.
Figure 6. Google Covid-19 open data (France).
1.3.3. Hidden algebraic equations.
Another problem is equation discovery. As an example let $w_{1},w_{2},w_{3},w_{4}$ and $x_{1},x_{2}$ be six variables such that the $w_{i}$ are i.i.d. $\mathcal{N}(0,1)$ random variables, and $x_{1},x_{2}$ satisfy $x_{1}=w_{1}w_{2}$ and $x_{2}=w_{2}\sin(w_{4})$. Then, given $N=1000$ observations of (samples from) those variables, consider the problem of discovering hidden functional relationships or dependencies between these variables (see Fig. 7.(a)).
Figure 7. Examples of Computational Hypergraph Discovery problem
1.3.4. Computational Biology.
An open problem in computational biology is to identify pathways and interactions between genes from their expression levels. One example stems from phosphorylated proteins and phospholipids in the human immune system T-cells presented in [34]. This dataset consists of $11$ proteins and $N=7446$ samples of their expressions. Fig. 7.(b) presents the (reference) causal dependencies found using experiments and structure learning algorithms for Bayesian networks, i.e., the directed acyclic graph identified in [34] from the cell-signaling data.
1.3.5. Land surface modeling.
Another example (see Fig. 7.(c) from [10]) is the modeling of land surface interactions in weather prediction, and the problem is to discover possibly hidden functional dependencies between state variables for a finite number of snapshots of those variables.
1.3.6. Economic network analysis.
Another example (see Fig. 7.(d) from [38]) is the recovery and analysis of economic networks, and the problem is to discover functional dependencies between the economic markers of different agents or companies, which is significant to systemic risk analysis.
1.3.7. Graph analysis of the stock market.
Another example (see Fig. 7.(e) from [44]) concerns a graph analysis of the stock market, and the problem is to discover functional dependencies between the price returns of different stocks.
1.3.8. Social network analysis.
Another example (see Fig. 7.(f) from [14]) is social network analysis/discovery, and the problem is to discover functional dependencies between quantitative markers associated with each individual (a vector of numbers per individual) in situations where the connectivity of the network may be hidden.
1.4. Objectives, contributions, and structure of this paper.
In this paper, we introduce an interpretable GP framework for Type 3 problems, targeting the data-driven discovery and completion of computational hypergraphs in the setting where the variables/nodes of the hypergraph are assumed to be known and its hyperedges/connectivity are unknown.
This work is organized as follows:
Sec. 2 presents a simplified overview of the proposed approach for Type 3 problems, with an application example, the Fermi-Pasta-Ulam-Tsingou (FPUT) problem.
Sec. 3 discusses the hardness of Type 3 problems (curse of combinatorial complexity, non-identifiability), proposes the well-posed formulation of such problems as a manifold learning problem, and reviews related causal inference approaches that could be employed for Type 3 problems under additional assumptions (access to the distributions of the underlying variables) which are not made here. In that sense, our method is different from all probabilistic approaches (targeting causal effects or conditional (in)dependencies).
Sec. 4 encapsulates the technical description of our method and presents the refinements needed to transform the simplified approach in Sec. 2 into a general-use algorithm. In particular, Sec. 4.1 shows that when the relations between variables are affine, then hypergraph discovery is equivalent to the Row Echelon Form reduction of the linear systems of equations encoded by the data. Sec. 4.2 presents a feature-map generalization of Row Echelon Form reduction from linear systems to nonlinear ones that can be represented by a given finite-dimensional feature-map given sufficient data. Sec. 4.3 presents a kernel-generalization of Row Echelon Form reduction encompassing arbitrarily complex functional dependencies between variables, low data regimes, regularization and generalization of free and dependent variables based on Kernel Mode Decomposition [30]. The proposed approach includes a kernel generalization of variance-based sensitivity analysis guiding the discovery of the structure of the hypergraph. Here, variables are linked via GPs, and those contributing to the highest data variance unveil the hypergraph’s structure. This GP variance decomposition of the data leads to signal-to-noise and a Z-score that can be employed to determine whether a given variable can be approximated as a nonlinear function of a given number of other variables.
Sec. 5 presents two algorithms/pseudocodes summarizing the proposed approach whose main steps are as follows: (1) For each variable of the hypergraph, try to approximate its values as a sum of noise and a function of the values of other variables in a given set of ancestors. (2) If the signal-to-ratio exceeds a given threshold, then iteratively remove variables that are least contributing to the signal until the signal-to-noise ratio drops a given threshold.
Sec. 6 provides a theoretical analysis of the signal-to-noise ratio which leads to a method for choosing the prior variance $\gamma$ of the noise.
Sec. 7 presents numerical experiments illustrating the scope and efficiency of the proposed approach. These include (algebraic) equation discovery, network discovery (gene pathways, chemical, and mechanical), and raw data analysis (Covid-19 open data).
@@ 1.3. Examples of Computational Hypergraph Discovery (Type 3) problems
2. Overview of the proposed approach for Type 3 problems
In this section, we present an algorithmic overview of the proposed approach for Type 3 problems. For ease of presentation, we consider the simple setting of Fig. 8.(a) where we are given $N$ samples on the variables $x_{1},\ldots,x_{6}$. After measurements/collection, these variables are normalized to have zero mean and unit variance. Our objective is to uncover the underlying dependencies between them. The code for this algorithm and its application to various examples are available for download (and as as an installable python library/package) in the Github repository of the paper.
2.1. A signal-to-noise ratio to decide whether or not a node has ancestors.
@@ 2. Overview of the proposed approach for Type 3 problems
To answer this question we regress $f$ with a centered GP $\xi\sim\mathcal{N}(0,\Gamma)$ whose covariance function $\Gamma$ is an additive kernel of the form
$\Gamma=K_{s}+\gamma\updelta(x-y)\,,$ (2.2)
where $K_{s}$ is a smoothing kernel, $\gamma>0$ and $\updelta(x-y)$ is the white noise covariance operator. This is equivalent to assuming the GP $\xi$ to be the sum of two independent GPs, i.e., $\xi=\xi_{s}+\xi_{n}$ where $\xi_{s}\sim\mathcal{N}(0,K_{s})$ is a smoothing/signal GP and $\xi_{n}\sim\mathcal{N}\big{(}0,\gamma\updelta(x-y)\big{)}$ is a noise GP. Writing $\mathcal{H}_{K_{s}}$ for the RKHS induced by the kernel $K$, this is also equivalent to approximating $f$ with a minimizer of
$\inf_{f\in\mathcal{H}_{K}}\|f\|_{K_{s}}^{2}+\frac{1}{\gamma}\big{\|}f(X)-Y\big{\|}_{\mathbb{R}^{N}}^{2}\,,$ (2.3)
where $\|\cdot\|_{\mathbb{R}^{N}}^{2}$ is the Euclidean norm on $\mathbb{R}^{N}$, $X$ is the input data on $f$ obtained as an $N$-vector whose entries $X_{i}$ are the samples on $x_{2},\ldots,x_{6}$, $Y$ is the output data on $f$ obtained as an $N$-vector whose entries are obtained from the samples on $x_{1}$, and $f(X)$ is a $N$-vector whose entries are the $f(X_{i})$. At the minimum
@@
quantifies the data variance explained by the noise GP $\xi_{n}$ [30]. This allows us to define the signal-to-noise ratio
@@
[…]
Figure 9. Selecting the signal kernel $K_{s}$.
2.2. Selecting the signal kernel $K_{s}$
This process is repeated by selecting the kernel $K_{s}$ to be linear ($K_{s}(x,x^{\prime})=1+\beta_{1}\sum_{i}x_{i}x_{i}^{\prime}$), quadratic ($K_{s}(x,x^{\prime})=1+\beta_{1}\sum_{i}x_{i}x_{i}^{\prime}+\beta_{2}\sum_{i\leq j}x_{i}x_{j}x_{i}^{\prime}x_{j}^{\prime}$) or fully nonlinear [^3] By fully nonlinear, we mean not linear and not quadratic. We also call such kernels interpolatory because we select them as having an infinite-dimensional feature map. These kernels are also known as universal kernels. Abusing language, we may also call a kernel interpolatory when the number of data points is smaller than the dimension of the feature map. to identify $f$ as linear, quadratic, or nonlinear. In the case of a nonlinear kernel, we employ:
@@ 2.2. Selecting the signal kernel $K_{s}$
where $k$ is a universal kernel, such as a Gaussian or a Matérn kernel, with all parameters set to $1$, and $\beta_{i}$ assigned the default value $0.1$. As illustrated in Fig. 9, we select $K_{s}$ as the first kernel that surpasses a signal-to-noise ratio of 0.5. If no kernel reaches this threshold, we conclude that $x_{1}$ lacks ancestors.
Figure 10. Pruning ancestors based on signal-to-noise ratio.
2.3. Pruning ancestors based on signal-to-noise ratio.
Once we establish that $x_{1}$ has ancestors, the next step is to prune its set of ancestors iteratively. We remove nodes with the least contribution to the signal-to-noise ratio and stop before that ratio drops below $0.5$ as illustrated in Fig. 10. To describe this, assume that $K_{s}$ is as in (2.7). Then $K_{s}$ is an additive kernel that can be decomposed into two parts:
@@
where $f_{1}$ does not depend on $x_{2}$, $f_{2}$ depends on $x_{2}$ and
$(f_{1},f_{2})=\operatorname{argmin}_{(g_{1},g_{2})\in\mathcal{H}_{K_{1}}\times\mathcal{H}_{K_{2}}}\|g_{1}\|_{K_{1}}^{2}+\|g_{2}\|_{K_{2}}^{2}\,.$ (2.10)
Furthermore,
$\|f\|_{K_{s}}^{2}=\|f_{1}\|_{K_{1}}^{2}+\|f_{2}\|_{K_{2}}^{2}\,,$ (2.11)
and $\frac{\|f_{2}\|_{K_{1}}^{2}}{\|f\|_{K_{s}}^{2}}\in[0,1]$ quantifies the contribution of $x_{2}$ to the signal data variance.

### v2

# Co-discovering Graphical Structure and Functional Relationships Within Data: A Gaussian Process Framework for Connecting the Dots

@@ Abstract.
Most problems within and beyond the scientific domain can be framed into one of the following three levels of complexity of function approximation. Type 1: Approximate an unknown function given input/output data. Type 2: Consider a collection of variables and functions, some of which are unknown, indexed by the nodes and hyperedges of a hypergraph (a generalized graph where edges can connect more than two vertices). Given partial observations of the variables of the hypergraph (satisfying the functional dependencies imposed by its structure), approximate all the unobserved variables and unknown functions. Type 3: Expanding on Type 2, if the hypergraph structure itself is unknown, use partial observations of the variables of the hypergraph to discover its structure and approximate its unknown functions. These hypergraphs offer a natural platform for organizing, communicating, and processing computational knowledge. While most scientific problems can be framed as the data-driven discovery of unknown functions in a computational hypergraph whose structure is known (Type 2), many require the data-driven discovery of the structure (connectivity) of the hypergraph itself (Type 3). We introduce an interpretable Gaussian Process (GP) framework for such (Type 3) problems that does not require randomization of the data, nor access to or control over its sampling, nor sparsity of the unknown functions in a known or learned basis. Its polynomial complexity, which contrasts sharply with the super-exponential complexity of causal inference methods, is enabled by the nonlinear analysis of variance capabilities of GPs used as a sensing mechanism.
Significance statement.
"Civilization advances by extending the number of important operations we can perform without thinking about them" (Whitehead, 1911). In line with this perspective, many complex data analysis problems within and beyond the scientific domain involve discovering graphical structures and functional relationships within data. Nonlinear variance decomposition with Gaussian Processes simplifies and automates this process. Other methods, such as Artificial Neural Networks, lack this variance decomposition feature. Information-theoretic and causal inference methods suffer from super-exponential complexity with respect to the number of variables. The proposed technique performs this task in polynomial complexity. This unlocks the potential for applications involving the identification of a network of hidden relationships between variables without a parameterized model at an unprecedented scale, scope, and complexity.
Introduction
The three levels of complexity of function approximation.
As illustrated in Fig. 1.(a-c), Type 1, Type 2 and Type 3 problems can be formulated as completing or discovering hypergraphs where nodes represent variables and edges represent functional dependencies. The graph in Type 1 has only two variables and one unknown function. The graph in Type 2 has multiple variables and (some possibly unknown) functions, and the connectivity of the graph is known. The graph in Type 3 has an unknown connectivity (functional dependencies between variables may be unknown) and this is the focus of this work. Current methods for solving Type 1 and 2 problems include Deep Learning (DL) methods, which benefit from extensive hardware and software support but have limited guarantees. Despite their prevalence, Type 3 challenges have been largely overlooked due to their inherent complexity. Causal inference methods morgan2015counterfactuals ; glymour2016causal and probabilistic graphs stegle2010probabilistic ; lopez2015towards and sparse regression methods doostan2011non ; brunton2016discovering , offer potential avenues for addressing Type 3 problems. However, it is important to note that their application to these problems necessitates additional assumptions. Causal inference models, for instance, typically assume randomized data and some level of access to the data generation process or its underlying distributions. Sparse regression methods, on the other hand, rely on the assumption that functional dependencies have a sparse representation within a known basis. In this paper, we do not impose these assumptions, and thus, these particular techniques may not be applicable. Furthermore while the complexity of Bayesian causal inference methods may grow super-exponentially with the number $d$ of variables, the complexity of our method is that of $d$ parallel computations of polynomial complexities bounded between $\mathcal{O}(d)$ (best case) and $\mathcal{O}(d^{4})$ (worst case).
@@ Introduction
Generalizing Gaussian Process methods.
Although Gaussian Process (GP) methods are sometimes perceived as a well-founded but old technology limited to curve fitting (Type 1 problems), they have recently been generalized, beyond Type 1 problems, to an interpretable framework (Computational Graph Completion or CGC owhadi2022computational ) for solving Type 2 problems owhadi2023ideas ; chen2021solving ; batlle2023error ; chen2023sparse ; darcy2023one ; hamzi2023learningpartial , all while maintaining the simple and transparent theoretical and computational guarantees of kernel/optimal recovery methods micchelli1977survey ; OwhScobook2018 . This paper introduces a comprehensive GP framework for solving Type 3 problems, which is interpretable and amenable to analysis. This framework leverages the Uncertainty Quantification (UQ) properties of GP methods, which do not have an immediate natural counterpart in DL methods. It is based on a kernel generalization wahba2003introduction ; owhadi2019kernelmd of variance-based sensitivity analysis guiding the discovery of the structure of the hypergraph. Here, variables are linked via GPs, and those contributing to the highest data variance unveil the hypergraph’s structure. This GP variance decomposition of the data leads to signal-to-noise and a Z-score that can be employed to determine whether a given variable can be approximated as a nonlinear function of a subset of other variables.
The scope of Type 1, 2 and 3 problems.
The scope of Type 1, 2 and 3 problems is immense. Numerical approximation OwhScobook2018 ; OwhScoSchNotAMS2019 ; SchaeferSullivanOwhadi17 ; schafer2021sparse , Supervised Learning, and Operator Learning KFforcimate2020 ; hamzi2021learning ; schafer2021sparsematrixvector ; batlle2023kernel can all be formulated as Type 1 problems, i.e., as approximating unknown functions given (possibly with noisy and infinite/high-dimensional) inputs/output data. The common GP based solution to these problems is to replace the underlying unknown function by a GP and compute its MAP estimator given available data. Type 2 problems include (Fig. 1.(d-h)) solving and learning (possibly stochastic) ordinary or partial differential equations chen2021solving ; darcy2023one , Deep Learning owhadi2023ideas , dimension reduction, reduced-ordered modeling, system identification owhadi2022computational , closure modeling, etc. Indeed, all these problems can be formulated as completing a computational graph owhadi2022computational . In this formulation, variables and functions are represented by the nodes and the edges of the graph whose structure corresponds to the functional dependencies between variables. Some of the functions and variables may be unknown, and by completing, we mean approximating the unknown functions (colored in red in Fig. 1) given samples from the observed variables. The common GP-based solution to Type 2 problems is to simply replace unknown functions by GPs and compute their MAP/MLE estimators given available data and constraints imposed by the structure of the graph owhadi2022computational . While most problems in Computational Sciences and Engineering (CSE) and Scientific Machine Learning (SciML) can be framed as Type 1 and Type 2 challenges, many problems in science can only be categorized as Type 3 problems, i.e., discovering the structure/connectivity of the graph itself from data prior to its completion. Indeed the scope of Type 3 problems extends well beyond Type 2 problems and includes equation discovery (Fig. 1.(i)); the modeling of land surface interactions in weather prediction (Fig. 1.(j) from dirmeyer2019land , discovering possibly hidden functional dependencies between state variables for a finite number of snapshots of those variables); social network analysis (Fig. 1.(k) from gittell2021relational , discovering functional dependencies between quantitative markers associated with each individual in situations where the connectivity of the network may be hidden); economic network analysis (Fig. 1.(m) from schweitzer2009economic , discovering functional dependencies between the economic markers of different agents or companies, which is significant to systemic risk analysis); and computational biology (Fig. 1.(l) from sachs2005causal , identifying pathways and interactions between genes from their expression levels).
@@
Overview of the proposed approach for Type 3 problems.
We first present an algorithmic overview of the proposed GP-based approach for Type 3 problems. For ease of presentation, we consider the simple setting of Fig. 2.(a) where we are given $N$ samples on the variables $x_{1},\ldots,x_{6}$. After measurements/collection, these variables are normalized to have zero mean and unit variance. Our objective is to uncover the underlying dependencies between them.
A signal-to-noise ratio to decide whether or not a node has ancestors.
@@
To answer this question we regress $f$ with a centered GP $\xi\sim\mathcal{N}(0,\Gamma)$ whose covariance function $\Gamma$ is an additive kernel of the form $\Gamma=K_{s}+\gamma\updelta(x-y)\,$, where $K_{s}$ is a smoothing kernel, $\gamma>0$ is a regularization parameter and $\updelta(x-y)$ is the white noise covariance operator. This is equivalent to assuming the GP $\xi$ to be the sum of two independent GPs, i.e., $\xi=\xi_{s}+\xi_{n}$ where $\xi_{s}\sim\mathcal{N}(0,K_{s})$ is a smoothing/signal GP and $\xi_{n}\sim\mathcal{N}\big(0,\gamma\updelta(x-y)\big)$ is a noise GP. Writing $\mathcal{H}_{K_{s}}$ for the Reproducing Kernel Hilbert Space (RKHS) induced by the kernel $K_{s}$, this is also equivalent to approximating $f$ with a minimizer of
$\inf_{f\in\mathcal{H}_{K_{s}}}\|f\|_{K_{s}}^{2}+\frac{1}{\gamma}\big\|f(X)-Y\big\|_{\mathbb{R}^{N}}^{2}\,$, (0.2)
where $\|\cdot\|_{\mathbb{R}^{N}}^{2}$ is the Euclidean norm on $\mathbb{R}^{N}$, $X$ is the input data on $f$ obtained as an $N\times 5$-matrix whose rows $X_{i}$ are the samples on $x_{2},\ldots,x_{6}$, $Y$ is the output data on $f$ obtained as an $N$-vector whose entries are obtained from the samples on $x_{1}$, and $f(X)$ is a $N$-vector whose entries are the evaluations $f(X_{i})$. At the minimum
@@
quantifies the data variance explained by the noise GP $\xi_{n}$ owhadi2019kernelmd . This allows us to define the signal-to-noise ratio
@@
If $\frac{\mathcal{V}(s)}{\mathcal{V}(s)+\mathcal{V}(n)}<0.5$ [^1] We will later present a version with a more sophisticated method for pruning, but we keep the 0.5 threshold in this example for simplicity., then, as illustrated in Fig. 2.(c), we deduce that $x_{1}$ has no ancestors, i.e., $x_{1}$ cannot be approximated as function of $x_{2},\ldots,x_{6}$. Conversely if $\frac{\mathcal{V}(s)}{\mathcal{V}(s)+\mathcal{V}(n)}>0.5$, then, we deduce that $x_{1}$ has ancestors, i.e., $x_{1}$ can be approximated as function of $x_{2},\ldots,x_{6}$.
Selecting the signal kernel $K_{s}$
This process is repeated by selecting the kernel $K_{s}$ to be linear ($K_{s}(x,x^{\prime})=1+\beta_{1}\sum_{i}x_{i}x_{i}^{\prime}$), quadratic ($K_{s}(x,x^{\prime})=1+\beta_{1}\sum_{i}x_{i}x_{i}^{\prime}+\beta_{2}\sum_{i\leq j}x_{i}x_{j}x_{i}^{\prime}x_{j}^{\prime}$) or fully nonlinear to identify $f$ as linear, quadratic, or nonlinear. In the case of a nonlinear kernel, we employ:
@@
where $k$ is a universal kernel, such as a Gaussian or a Matérn kernel, with all parameters set to $1$, and $\beta_{i}$ assigned the default value $0.1$. We select $K_{s}$ as the first kernel that surpasses a signal-to-noise ratio of 0.5. If no kernel reaches this threshold, we conclude that $x_{1}$ lacks ancestors.
Pruning ancestors based on signal-to-noise ratio.
Once we establish that $x_{1}$ has ancestors, the next step is to prune its set of ancestors iteratively. We remove nodes with the least contribution to the signal-to-noise ratio and stop before that ratio drops below $0.5$ as illustrated in Fig. 2.(d). To describe this, assume that $K_{s}$ is as in (0.6). Then $K_{s}$ is an additive kernel that can be decomposed into two parts: